# A Reduction-Driven Local Search for the Generalized Independent Set Problem

### v1

@@ B Comparative Details on 216 Instances of SET-1
brock200_2_A 200 2438 7438 2999 10800 3078 10800 3326 8 3326 $<$1 3326 $<$1
brock400_2_A 400 14751 45035 2603 10800 2787 10800 2941 12 2941 1 2941 3
C125.9_A 125 1733 5230 1837 365 1837 62.71 1837 $<$1 1837 $<$1 1837 $<$1
C250.9_A 250 7073 20911 2011 10800 2171 10800 2171 5 2171 $<$1 2171 $<$1
gen200_p0.9_55_A 200 4425 13485 2094 10800 2096 7832 2096 $<$1 2096 $<$1 2096 $<$1
gen400_p0.9_75_A 400 18063 53757 2164 10800 2310 10800 2404 19 2404 $<$1 2404 1
hamming8-4_A 256 5173 15691 2813 10800 3124 10800 3124 1 3124 $<$1 3124 $<$1
keller4_A 171 2400 7035 2532 10800 2690 10800 2690 $<$1 2690 $<$1 2690 $<$1
MANN_a27_A 378 17580 52971 1976 10800 2206 10800 2208 $<$1 2208 $<$1 2208 $<$1
p_hat300-1_A 300 2734 8199 5566 10800 $6$ 10800 7899 25 7899 $<$1 7899 $<$1
p_hat300-2_A 300 5603 16325 4102 10800 4714 10800 5343 $<$1 5343 $<$1 5343 $<$1
p_hat300-3_A 300 8388 25002 2477 10800 2658 10800 2838 2 2838 $<$1 2838 $<$1
brock200_2_B 200 2438 7438 1398 10800 1427 10800 1533 $<$1 1533 $<$1 1533 $<$1
brock400_2_B 400 14751 45035 1161 10800 1210 10800 1291 16 1291 1 1291 4
C125.9_B 125 1733 5230 837 523 837 84 837 $<$1 837 $<$1 837 $<$1
C250.9_B 250 7073 20911 911 10800 969 10800 971 3 971 $<$1 971 $<$1
gen200_p0.9_55_B 200 4425 13485 944 10800 946 10800 946 $<$1 946 $<$1 946 $<$1
gen400_p0.9_75_B 400 18063 53757 936 10800 1010 10800 1052 11 1054 1 1054 1
hamming8-4_B 256 5173 15691 1309 10800 1474 10800 1474 $<$1 1474 $<$1 1474 $<$1
keller4_B 171 2400 7035 1190 10800 1240 10800 1254 $<$1 1254 $<$1 1254 $<$1
MANN_a27_B 378 17580 52971 875 10800 956 10800 958 11 958 $<$1 958 $<$1
p_hat300-1_B 300 2734 8199 2666 10800 3216 10800 3818 2 3818 $<$1 3818 $<$1
p_hat300-2_B 300 5603 16325 1952 10800 2177 10800 2543 $<$1 2543 $<$1 2543 $<$1
p_hat300-3_B 300 8388 25002 1101 10800 1208 10800 1288 2 1288 $<$1 1288 $<$1
brock200_2_C 200 2438 7438 607 10800 607 10800 658 $<$1 658 $<$1 658 $<$1
brock400_2_C 400 14751 45035 461 10800 465 10800 500 3 500 $<$1 500 $<$1
C125.9_C 125 1733 5230 337 1519 337 204 337 $<$1 337 $<$1 337 $<$1
C250.9_C 250 7073 20911 359 10800 370 10800 372 5 372 $<$1 372 $<$1
gen200_p0.9_55_C 200 4425 13485 352 10800 371 10800 371 $<$1 371 $<$1 371 $<$1
gen400_p0.9_75_C 400 18063 53757 356 10800 379 10800 395 16 395 $<$1 395 $<$1
hamming8-4_C 256 5173 15691 556 10800 649 10800 652 $<$1 652 $<$1 652 $<$1
keller4_C 171 2400 7035 510 10800 554 10800 558 $<$1 558 $<$1 558 $<$1
MANN_a27_C 378 17580 52971 325 10800 335 10800 335 11 335 $<$1 335 $<$1
p_hat300-1_C 300 2734 8199 1205 10800 1491 10800 1793 1 1793 $<$1 1793 $<$1
p_hat300-2_C 300 5603 16325 854 10800 981 10800 1159 $<$1 1159 $<$1 1159 $<$1
p_hat300-3_C 300 8388 25002 453 10800 494 10800 529 1 529 $<$1 529 $<$1

### v2

@@ B Comparative Details on 216 Instances of SET-1
brock200_2_A 200 2,438 7,438 2,999 10,800 3,078 10,800 3,326 8 3,326 <<1 3,326 <<1
brock400_2_A 400 14,751 45,035 2,603 10,800 2,787 10,800 2,941 12 2,941 1 2,941 3
C125.9_A 125 1,733 5,230 1,837 365 1,837 62.71 1,837 <<1 1,837 <<1 1,837 <<1
C250.9_A 250 7,073 20,911 2,011 10,800 2,171 10,800 2,171 5 2,171 <<1 2,171 <<1
gen200_p0.9_55_A 200 4,425 13,485 2,094 10,800 2,096 7,832 2,096 <<1 2,096 <<1 2,096 <<1
gen400_p0.9_75_A 400 18,063 53,757 2,164 10,800 2,310 10,800 2,404 19 2,404 <<1 2,404 1
hamming8-4_A 256 5,173 15,691 2,813 10,800 3,124 10,800 3,124 1 3,124 <<1 3,124 <<1
keller4_A 171 2,400 7,035 2,532 10,800 2,690 10,800 2,690 <<1 2,690 <<1 2,690 <<1
MANN_a27_A 378 17,580 52,971 1,976 10,800 2,206 10,800 2,208 <<1 2,208 <<1 2,208 <<1
p_hat300-1_A 300 2,734 8,199 5,566 10,800 6,666 10,800 7,899 25 7,899 <<1 7,899 <<1
p_hat300-2_A 300 5,603 16,325 4,102 10,800 4,714 10,800 5,343 <<1 5,343 <<1 5,343 <<1
p_hat300-3_A 300 8,388 25,002 2,477 10,800 2,658 10,800 2,838 2 2,838 <<1 2,838 <<1
brock200_2_B 200 2,438 7,438 1,398 10,800 1,427 10,800 1,533 <<1 1,533 <<1 1,533 <<1
brock400_2_B 400 14,751 45,035 1,161 10,800 1,210 10,800 1,291 16 1,291 1 1,291 4
C125.9_B 125 1,733 5,230 837 523 837 84 837 <<1 837 <<1 837 <<1
C250.9_B 250 7,073 20,911 911 10,800 969 10,800 971 3 971 <<1 971 <<1
gen200_p0.9_55_B 200 4,425 13,485 944 10,800 946 10,800 946 <<1 946 <<1 946 <<1
gen400_p0.9_75_B 400 18,063 53,757 936 10,800 1,010 10,800 1,052 11 1,054 1 1,054 1
hamming8-4_B 256 5,173 15,691 1,309 10,800 1,474 10,800 1,474 <<1 1,474 <<1 1,474 <<1
keller4_B 171 2,400 7,035 1,190 10,800 1,240 10,800 1,254 <<1 1,254 <<1 1,254 <<1
MANN_a27_B 378 17,580 52,971 875 10,800 956 10,800 958 11 958 <<1 958 <<1
p_hat300-1_B 300 2,734 8,199 2,666 10,800 3,216 10,800 3,818 2 3,818 <<1 3,818 <<1
p_hat300-2_B 300 5,603 16,325 1,952 10,800 2,177 10,800 2,543 <<1 2,543 <<1 2,543 <<1
p_hat300-3_B 300 8,388 25,002 1,101 10,800 1,208 10,800 1,288 2 1,288 <<1 1,288 <<1
brock200_2_C 200 2,438 7,438 607 10,800 607 10,800 658 <<1 658 <<1 658 <<1
brock400_2_C 400 14,751 45,035 461 10,800 465 10,800 500 3 500 <<1 500 <<1
C125.9_C 125 1,733 5,230 337 1,519 337 204 337 <<1 337 <<1 337 <<1
C250.9_C 250 7,073 20,911 359 10,800 370 10,800 372 5 372 <<1 372 <<1
gen200_p0.9_55_C 200 4,425 13,485 352 10,800 371 10,800 371 <<1 371 <<1 371 <<1
gen400_p0.9_75_C 400 18,063 53,757 356 10,800 379 10,800 395 16 395 <<1 395 <<1
hamming8-4_C 256 5,173 15,691 556 10,800 649 10,800 652 <<1 652 <<1 652 <<1
keller4_C 171 2,400 7,035 510 10,800 554 10,800 558 <<1 558 <<1 558 <<1
MANN_a27_C 378 17,580 52,971 325 10,800 335 10,800 335 11 335 <<1 335 <<1
p_hat300-1_C 300 2,734 8,199 1,205 10,800 1,491 10,800 1,793 1 1,793 <<1 1,793 <<1
p_hat300-2_C 300 5,603 16,325 854 10,800 981 10,800 1,159 <<1 1,159 <<1 1,159 <<1
p_hat300-3_C 300 8,388 25,002 453 10,800 494 10,800 529 1 529 <<1 529 <<1